40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 319 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  929   Thu Sep 4 17:44:27 2008 YoichiUpdatePSLFSS error signal spectrum
Attached is a spectrum of the FSS error signal.
There are a lot of sharp peaks above 100kHz (the UGF of the servo is about 200kHz).
These are mostly harmonics of 77kHz. They are the current suspects of the FSS slew rate saturation.
I remember when I blocked the light to the PD, these peak went away. So these noises must be
in the light. But I checked it a few weeks ago. So I will re-check it later.

One possible source of the lines is a DC-DC converter in the NPRO near the crystal.
We will try to move the converter outside of the box.
Attachment 1: FSS-Error-Spe.png
FSS-Error-Spe.png
  928   Thu Sep 4 17:17:03 2008 YoichiUpdateIOOMC open loop TF
I measured open loop transfer functions of the MC servo.
The UGF was about 30kHz. Since there was some gain margin at higher frequencies, I increased
the input gain of the MC servo board from 19dB to 22dB. Now the UGF is 40kHz and we have more
phase margin (~30deg).
Attachment 1: MC-OPLTF.png
MC-OPLTF.png
  927   Thu Sep 4 17:12:57 2008 YoichiUpdatePSLFSS open loop TF
I changed the gain settings of the FSS servo.
Now the Common Gain is 5dB (the last night it was 2dB) and the Fast Gain is 12dB (formerly 16dB).
I measured the open loop TF with this setting (the attachment).
I also plotted the OPLTF when CG=2dB, FG=20.5dB. With this setting, the MC looses lock every 30min.

You can see that the OPLTF is smoother with FG=12dB.
When the FG is high, you can see some structure around 250kHz. This structure is reproducible.
This may be some interruption from the fast path to the PC path through a spurious coupling.
Attachment 1: FSS-OPLTFs.png
FSS-OPLTFs.png
  926   Thu Sep 4 17:03:25 2008 YoichiUpdatePSLRF oscillator noise comparison
I measured current spectra of the RF signal going to the FSS EOM.
The attachment compares the spectra between a Stanford signal generator and a Marconi.
I borrowed the Marconi from the abs. length measurement experiment temporarily.
The measurement was done using the signal going to the EOM. That means the spectra include
noise contributions from the RF amp., splitter and cables.

21.5MHz peak was not included because that would overload the ADC and I would have to use a large attenuation.
This means the measurement would be totally limited by ADC noise everywhere except for 21.5MHz.

I noticed that with the Marconi, the FSS is a little bit happier, i.e. the PC path is less loaded
(0.9Vrms with Stanford vs. 0.7Vrms with Marconi). But the difference is small.
Probably the contribution from the 77kHz harmonics in the laser light is more significant (see entry #929).
Also the peaks in the Stanford spectrum are not harmonics of 77kHz, which we see in the FSS error signal.

I returned the Marconi after the measurement to let Alberto work on the abs. length measurement.
Attachment 1: RFSpectra.png
RFSpectra.png
  925   Thu Sep 4 16:24:56 2008 ranaConfigurationComputersAttempt to increase gain for C1:PSL-ISS_INMONPD_F via 110B

Quote:
We were attempting to increase the gain on the channel C1:PSL-ISS_INMONPD_F in preparation to do a scan of the PMC at very low input power.

According to the Wikipedia, certain esoteric mathematical
operations lead to the result that 4000 x 10 > 32768.
  924   Thu Sep 4 14:43:58 2008 JenneUpdatePSLPMC Open Loop Gain
I have measured the PMC's open loop gain. UGF is 629.7Hz, with a phase margin of 53 degrees.

I injected into FP2 on the front panel, and measured MixOut/Source from 100Hz to 100kHz using the SR785. I did this both when the loop was open, and when the loop was closed (open the loop by enabling FP1, which breaks the loop).

We have 2 transfer functions involved: The actual open loop gain of the PMC servo loop (G1), and the gain between FP2 and the MixerOut monitor point (G2). This gives us:

TF(closed loop) = G2*(1+G1)
TF(broken loop) = G2

G1 = TF(closed)/TF(broken) - 1

This G1 is the final open loop gain, and it is plotted below.
Attachment 1: OpenLoopTF04Sept2008.png
OpenLoopTF04Sept2008.png
  923   Thu Sep 4 13:48:50 2008 YoichiUpdatePSLFSS modulation depth
I scanned the reference cavity with the NPRO temperature (see the attached plot).
The power ratio between the carrier and the sideband resonances is about 26.8.
It corresponds to gamma=0.38.
The RF power fed into the EOM is now 14.75dBm (i.e. 1.7V amplitude). The NewFocus catalog says 0.1-0.3rad/V. So
gamma=0.38 is a reasonable number.




Attachment 1: RCScan.png
RCScan.png
  922   Thu Sep 4 11:33:25 2008 josephb, Eric, JenneConfigurationComputersAttempt to increase gain for C1:PSL-ISS_INMONPD_F via 110B
We were attempting to increase the gain on the channel C1:PSL-ISS_INMONPD_F in preparation to do a scan of the PMC at very low input power.

We started by adding a line to the C1:IOOF.ini file in /cvs/cds/caltech/chans/daq/ under that channel that said "gain=10.0". Before touching anything, the channel was outputting around 4000 counts.

We hit the reconfig button for c1iovme16k, then rebooted c1iovme (which turned out to do nothing) and then the framebuilder, in a method consistent with the wiki. This turned out to put the channel in an odd state, where it was showing very rapid, random spikes, virtually but still around 4000ish counts. We returned the file back to its original format, hit reconfig, and then rebooted the framebuilder. The channel however, was still behaving in the same broken way.

After poking around the PSL table, looking at some direct outputs, we came back and rebooted c1iovme and the framebuilder again, which fixed the channel, such that it was reading out correctly. Taking this as a sign that maybe we should reboot the framebuilder, then c1iovme to get the channel to load changes, we changed the file again to have "gain=10.0". Upon reboot of the framebuilder, the channel was still reading out fine, but at the same level. So we continued with the reboot of c1iovme. This still had no effect on the channel output.

The ini file has been set back at this point, however since Yoichi is working, I'm holding off doing a reconfig and reboot on the framebuilder until later.
  921   Thu Sep 4 10:13:48 2008 JenneUpdateIOOWe unlocked the MC temporarily
[Joe, Eric, Jenne]

While trying to diagnose some DAQ/PD problems (look for Joe and Eric's entry later), we unlocked the PMC, which caused (of course) the MC to unlock. So if you're looking back in the data, the unlock at ~10:08am is caused by us, not whatever problems may have been going on with the FSS. It is now locked again, and looking good.
  920   Thu Sep 4 07:46:10 2008 YoichiUpdateIOOMC is now happy
The MC has been locked for more than 12 hours continuously now !
Changes I made yesterday were:
(1) Removed the 20dB attenuator before the EOM.
(2) Reduced the Fast Gain from 21dB to 16dB, which made the PC to be a little bit more loaded (~0.6Vrms).

As Rana pointed out in the meeting, setting the Fast Gain a bit lower may have put the FSS in a stabler state.
Attachment 1: MC-lock.png
MC-lock.png
  919   Thu Sep 4 07:29:52 2008 YoichiUpdatePSLc1iovme power cycled

Quote:
Entry 663 has a plot of this using the PSL/FSS/SLOWscan script. It shows that the SB's were ~8x smaller than the carrier.
P_carrier   J_0(Gamma)^2 
--------- = ------------
P_SB        J_1(Gamma)^2

Which I guess we have to solve numerically for large Gamma?


P_carrier/P_SB = 8 yields gamma=0.67.
  918   Thu Sep 4 00:38:14 2008 ranaUpdatePSLc1iovme power cycled
Entry 663 has a plot of this using the PSL/FSS/SLOWscan script. It shows that the SB's were ~8x smaller than the carrier.
P_carrier   J_0(Gamma)^2 
--------- = ------------
P_SB        J_1(Gamma)^2

Which I guess we have to solve numerically for large Gamma?
  917   Wed Sep 3 19:09:56 2008 YoichiDAQComputersc1iovme power cycled
When I tried to measure the sideband power of the FSS using the scan of the reference cavity, I noticed that the RC trans. PD signal was not
properly recorded by the frame builder.
Joe restarted c1iovme software wise. The medm screen said c1iovme is running fine, and actually some values were recorded by the FB.
Nonetheless, I couldn't see flashes of the RC when I scanned the laser frequency.
I ended up power cycling the c1iovme and run the restart script again. Now the signals recorded by c1iovme look fine.
Probably, the DAQ boards were not properly initialized only by the software reset.
I will re-try the sideband measurement tomorrow morning.
  916   Wed Sep 3 18:45:01 2008 AlbertoConfiguration PD3 gain
Alberto, Yoichi,

We found that the PD3 servo was unstable with a gain of 1, so we switched it to 0.5
  915   Wed Sep 3 18:43:19 2008 YoichiConfigurationElectronicsTwo more active probes
I found two active probes, an HP41800A and a Tektronix P6201.
Thanks John for telling me you saw them before.
Now we have three active probes, wow !
We have to find or make a power supply for P6201.
The manual of the probe is attached.
Attachment 1: Tektronix-P6201-active-probe.pdf
Tektronix-P6201-active-probe.pdf Tektronix-P6201-active-probe.pdf Tektronix-P6201-active-probe.pdf Tektronix-P6201-active-probe.pdf Tektronix-P6201-active-probe.pdf Tektronix-P6201-active-probe.pdf Tektronix-P6201-active-probe.pdf Tektronix-P6201-active-probe.pdf
  914   Wed Sep 3 12:26:49 2008 EricSummaryCamerasWeekly Summary
Finished up simulating the end mirror error in order to test the whether the fitting code still provides reasonable answers despite the noise caused by the defects on the end mirror. The model I used to simulate the defects is far from perfect, but its good enough given the time I have remaining, and I have no reason to believe the differences between it and the real noise would cause any radical changes in how the fit operates. A comparison between a modeled image and a real image is attached. Average error (difference between the estimated value and the real value) for each of the parameters is

For the fit:
Max Intensity: 2767.4 (Max intensities ranged from 8000 to 11000)
X-Position: 0.9401 pixels
X Beam Waist: 1.3406 pixels (beam waists ranged from 35 to 45)
Y-Position: 0.9997 pixels
Y Beam Waist: 1.3059 pixels (beam waists ranged from 35 to 45)
Intensity Offset: 12.7705 (Offsets ranged from 1000 to 4000)

For the center of mass calculation (with a threshold that cut off everything above 13000)
X-Position: 0.0087 pixels
Y-Position: 0.0286 pixels

Thus, the fit is generally trustworthy for all parameters except for maximum intensity, for which it is very inaccurate. Additionally, this shows that the center of mass calculation actually does a much better job than the fit when this much noise is in the image. For the end mirrors, the fit is really only useful for finding beam waist, and even this is not extremely accurate (~3% error). All the parameters for the modeling is on the svn in /trunk/docs/emintun/MatLabFiles/EndMirrorErrorSimulation.txt.

Finished working on the calculations that convert a beam misalignment as measured as a change in the beam position on the two mirrors to a power loss in the cavity. Joe calculated the minimum measurable change in beam position to be around a tenth of a pixel, which corresponds to half a micron when the beam is directly incident on the camera. This gives the ability to measure fractional power losses as low as 2*10^-10 for the 40m main arm cavities. To me, this seems unusually low, though it scales with beam position squared, so if anything else limited the ability to measure changes in the beam position, it would have a large effect on the sensitivity to power losses. Additionally, it scales inversely with length, so shorter cavities provide less sensitivity.

This morning Joe and I tested the ability for the camera code to servo the ITMX in order to change the beam's position on the ETMX. Two major things have been changed since the last time we tried this. First, the calculated beam center that gets output to the EPICS channels now first goes through a transform that converts it from pixels into physical units, and should account for the oblique angle of the camera. The output to the EPICS channels should now be in the form of 'mm from the center of the optic', although this is not very precise at the moment. The second thing that was changed was that the servo was run with a modified servo script that included options to set a minimum, maximum, and slew rate in order to protect the mirrors from being swung around too much. The servo was generally successful: for a given x-position, it was capable of changing the yaw of ITMX so that the position seen on the camera moved to this new location. The biggest problem is that the x and y dimensions do not appear to be decoupled (the transform converting it to physical units should have done this), so that modifying the yaw of the mirror changed both the x and y positions (the y about half as much) as output by the camera. This could cause a problem when trying to servo in both dimensions at once, since one servo could end up opposing the other. I don't know the cause of this problem yet, since the transform that is currently in use appears to be correctly orienting the image.
Attachment 1: SimulatedErrorComparison.png
SimulatedErrorComparison.png
  913   Tue Sep 2 22:43:16 2008 YoichiConfigurationPSLUpdated FSS open loop TF
Since the LO level of the FSS servo was too low, I replaced the RF oscillator board with a combination of
a Stanford signal generator and an RF amplifier.
Right now, the POY RF amplifier is used for this purpose temporarily.
Now the LO level is about 16dBm. The RF power going into the EOM is attenuated by 20dB from the LO level.
I played with the cable length to get the phase right.
Then I was able to lock the FSS with the new RF signal source.

Attached is the open loop transfer function of the current FSS. Now the UGF is a bit above 200kHz, a factor of 2 improvement.
This gain was achieved with the common gain slider at 13.5dB and the fast gain = 30dB.
With the old RF oscillator board, UGF=100kHz was achieved with the common gain =30dB. Therefore, the increase of the LO gave
us a large signal gain.

Increasing the gain further, again ,makes the PC path crazy.
Rich suggested that this craziness was caused either by the slew rate limit of the PA85 or the output voltage limit of the bypass Op-amp(A829)
is hit.

TO DO:
* Look at the error signal spectrum to see if there is any signal causing the slew rate saturation at high frequencies.
* Find out what the RF signal level for the EOM should be. 20dB attenuation is an arbitrary choice.
* Find out the cross over frequency. Determine where the fast gain slider should be.
etc ...
Attachment 1: OPLTF.png
OPLTF.png
  912   Tue Sep 2 14:28:41 2008 YoichiUpdatePSLFSS EOM driving signal spectra
Rich advised me to change the +10V input of the FSS crystal frequency reference board from whatever voltage supply we use now to a nice one.
This voltage is directory connected to the signal lines of both LO and RF output amps. Therefore, fluctuations in the voltage directly appear
in the outputs, though DC components are cut off by the AC coupling capacitors.

I changed the source of this voltage from the existing Sorensen one to a power supply sitting next to the rack.
The attached plots shows the difference of the RF output spectra between the two 10V sources.
The low frequency crap is almost gone in the new 10V spectrum.

I tried to increase the FSS gain with the new 10V, but still it goes crazy. I suspect it is because the LO power is too low.
Attachment 1: RFDrive1.png
RFDrive1.png
Attachment 2: RFDrive2.png
RFDrive2.png
  911   Tue Sep 2 10:09:03 2008 steveUpdatePSLhead temp is cooling down
The chiller was over flowing this morning.
800 cc of water was removed.
PSL-126MOPA_HTEMP peaked at 20.7 C (normal is 18.7 C)
  910   Tue Sep 2 09:58:42 2008 YoichiConfigurationPSLFSS on an auxiliary loop

Quote:
Summary: The FSS is now temporarily disabled. Naturally, the MC won't lock. I will fix it tomorrow morning.


Now I removed the power splitters for the aux. reference cavity servo. The FSS is back and the MC locks.
I'm now returning one of the active high-impedance probes to the Wilson house. They need it today.
We are left with only one active probe. If anyone finds another active probe in the 40m lab.,
please let me know (according to Rana we should have one more).
  909   Tue Sep 2 07:58:34 2008 ranaSummaryPSLFSS & PMC LO trends for 2 years
The attached plot is a 2 year minute trend of the EPICS readback of the PMC & FSS LO Monitors (FSS_LODET & PMC_LODET).
Clearly the FSS LO has been dying for at least 2 years. The step up from 10 months
ago is probably when Rob removed a 3dB attenuator from in front of the box.
Attachment 1: psl-lo-trend.png
psl-lo-trend.png
  908   Mon Sep 1 19:23:17 2008 YoichiConfigurationPSLFSS on an auxiliary loop
Summary: The FSS is now temporarily disabled. Naturally, the MC won't lock. I will fix it tomorrow morning.

Today, I did the 4th item of my TO DO list.
Using a mini-circuit mixer and two SR560s, I constructed an auxiliary servo loop for the reference cavity.
With this loop, I was able to lock the reference cavity without using the FSS box.
By locking the reference cavity with this auxiliary servo, I was able to measure the PC path transfer function.
I will post the analyzed results later.

I borrowed the PD RF and the LO signals from the main FSS loop by power splitters. Therefore, the gain of the main FSS loop
is now about 3dB low. I tried to compensate it by increasing the EOM modulation depth, but the PC path is still a bit noisy.
Probably the already too low LO power is now seriously low (the LO power cannot be changed from EPICS).
Because I did not want to leave the PC path with large output overnight (it will heat up the PA85, and might cause damage, though unlikely),
I disabled the FSS for now.
  907   Mon Sep 1 04:34:00 2008 ranaUpdatePSLFSS loop transfer functions
I started from 6th item in Yoichi's todo list.

1) Increased the setpoint of the thermostat next to the framebuilder from 73F to 79F. Its freezing over there
in the room with the drill press. Steve's illegal mercury thermometer is reading 19 C.


2) Looked the RFPD's output spectrum using the 20 dB coupled output from the coupler that's in-line.
The first attached PDF file (n.pdf) has several plots:
page 1: 0-500 MHz anomolous peaks at 138 & 181 MHz but nothing too crazy
page 2: 0-100 MHz 80 MHz peak is RF pickup from the VCO Driver - not on the light
page 3: 10-30 MHz totally nuts
page 4: 18-25 MHz that's just wrong

The RF spectrum should only have some action around 21.5 MHz and a little peak at 2x 21.5 MHz. All that extra
junk means that something is broken!


3) To see if I could rid of any of the 80 MHz signal or any of that other trash from 18-25 MHz, I wound the RF cable
around a large toroidal ferrite core. This should have given us many uH of inductance for any signals common to
both the center and shield of the cable with no effect on the differential RF signals. There was no effect.


4) Next went to look at the 21.5 MHz Crystal Oscillator Reference card (D980353...I bet you can't figure out how
this one works). These have the Mini-Circuits SMA 30 MHz low pass (SLP-30) filters on both the LO and EOM outputs.

FSSLO.PNG shows the waveform after 20 dB attenuation going into a scope terminated with 50 Ohms.
FSSLO-Spec.png shows the spectrum of this signal - its pretty distorted. Here's the levels
   f (MHz) |  before filter (dBm) | after filter (dBm)
   ---------------------------------------------------
     21.5  |       -12.8               -13.1       
     43            -24                 -46
     64.5          -50               < -80
     86            -64               < -80

This would be OK after the filter, but the level is very low. Only 7dBm (accounting for my 20 dB att) !!
The FSS uses a JMS-1H mixer which needs, as everyone knows, a +17 dBm LO signal. Que lastima.

There seems to be something wrong already, but wait...


5) PC25.PNG shows the output signal going to the EOM from 0 - 25 MHz. The step that's visible there at
around 10 MHz is just something inherent to the analyzer (??). But see all that crap there down below
5 MHz ? That is NOT supposed to be there.

pc.pdf shows on the first page the comparison in EOM drive with 2 different slider values on the
RF AM adjust screen for the FSS. But page 2 is the punchline of this long entry: There is a bunch of
excess junk on the drive signal going to the FSS's phase modulator.
The FSS is then trying to handle
this extra frequency noise and getting into trouble.

We have to fix this board. I have also ordered a few SBP-21.4 from mini-circuits (SMA bandpass around 21.4 MHz)
just in case. Another option is to just replace this thing with a Marconi and an RF amp.






Attachment 1: n.pdf
n.pdf n.pdf n.pdf n.pdf
Attachment 2: FSSLO.PNG
FSSLO.PNG
Attachment 3: FSSLO-Spec.png
FSSLO-Spec.png
Attachment 4: PC25.png
PC25.png
Attachment 5: pc.pdf
pc.pdf pc.pdf
  906   Sat Aug 30 13:28:01 2008 ranaConfigurationPSLPMC: List of changes
This is a list of changes made to the PMC board while we had it out for modifying the notch:

  • LC-LC 4th order low pass filter
  • Replace the AD797 (U2) with an OP27. AD797's are bad - do not use them anywhere for any reason. The OP27 is slower and has a 3x worse input noise but doesn't compromise the bandwidth or noise performance of the PMC by any significant amount. The rule is: use OP27 everywhere unless you have a very good reason why not.
  • There is no 'H1' jumper on board. R9 is 90.9 Ohms and R2=900 Ohms so that the U2 stage has a gain of 10.
  • Cut a trace and inserted a 500 Ohm resistor between U2-pin6 and U5A-pin2 (the AD602). The AD602 has a 100 Ohm input impedance which cannot be driven without limiting by the AD797 or the OP27. The 500 Ohm resistor makes it a driveable load for low level signals which is all that should be there since its the error point of the servo. it also becomes a 6:1 voltage divider. Since the AD602 has a fixed output voltage noise of 100 nV/rHz, this will limit the noise performance if the VGA gain is less than 20 dB, but whatever.
  • R11 7.87k -> 1.74k, R12 = 78.7k -> 700k. This increases the high frequency gain of that stage by 7.87/1.74 = 4.5 and lowers the low frequency pole from 2 to 0.2 Hz to give the PMC some more staying power at DC. The loop shape is now 1/f^2 in the 9-480 Hz band and so the phase dips enough to make it almost conditionally stable, but not quite.
  • C26 changed from ??? + a 30 pF trim cap into a fixed NNN pF cap to set the notch frequency for the 14.5 kHz body mode that we measured. Once our brick configuration is more settled we can increase the Q of this notch from small to big.
  • Grounded pin 5 of U14 & U15 (AD620). These have sometimes been used as "differential" drivers in LIGO by connecting this reference voltage pin to the remote ground of the next board. This has always lead to insidious oscillation and noise. This beauty also has an output noise of 100 nV/rHz. Just never use this chip if you can help it; we can make true differential drivers - we have the technology.

Of course, we didn't have a current version of a schematic sitting around so I printed out a Rev E schematic and marked it up with red pen. I'll post pictures later and put the schematic into the PSL schematics notebook. Would be useful to take the old schematic and update it in Acrobat so that we have something electronic.
  905   Fri Aug 29 22:57:48 2008 YoichiUpdatePSLFSS loop transfer functions
I've been measuring a bunch of transfer functions of the FSS related stuffs.
There are a lot to be analyzed yet, but here I put one mystery I'm having now.
Maybe I'm missing something stupid, so your suggestions are welcome.

Here is a conceptual diagram of the FSS control board

                                                          TP3             TP4
                                                           ^               ^
                                                           |               |
RF PD -->--[Mixer]-----[Sum Amp]------>--[Common Gain]--->----[Fast Gain]----[Filter]--> NPRO PZT
              ^     |      ^        |                  |     
              |     V      |        V                  |
LO ---->-------    TP1     IN      TP2                 -->---[Filter]--[High Volt. Amp.] --> Phase Corrector

What I did was first to measure a "normal" openloop transfer function of the FSS servo.
The FSS was operated in the normal gain settings, and a signal was injected from "IN" port.
The open loop gain was measured by TP1/TP2.
Now, I disconnected the BNC cable going to the phase corrector to disable the PC path and locked the ref. cav. 
only using the PZT. This was done by reducing the "Common Gain" and "Fast Gain" by some 80dB.
Then I measured the open loop gain of this configuration. The UGF in this case was about 10kHz.
I also measured the gain difference between the "normal" and "PZT only" configurations by injecting 
a signal from "IN" and measuring TP3/TP2 and TP4/TP3 with both configurations (The signal from the Mixer was
disconnected in this measurement). 

The first attachment shows the normal open loop gain (purple) and the PZT only open loop gain scaled by the 
gain difference (about 80dB). The scaled PZT open loop gain should represent the open loop gain of the PZT
path in the normal configuration. So I expected that, at low frequencies, the scaled PZT loop TF overlaps the normal
open loop TF.
However, it is actually much larger than the normal open loop gain.
When I scale the PZT only TF by -30dB, it looks like the attachment #2.
The PZT loop gain and the total open loop gain match nicely between 20kHz and 70kHz.
Closer look will show you that small structures (e.g. around 30kHz and 200kHz) of the two
TFs also overlap very well. I repeated measurements many times and those small structures are always there (the phase is
also consistently the same). So these are not random noise.

I don't know where this 30dB discrepancy comes from. Is it the PC path eating the PZT gain ?

I have measured many other TFs. I'm analyzing these.
Here is the TO DO list:

* Cavity response plot from AOM excitation measurements.
* Cavity optical gain plot.
* Reconstruct the open loop gain from the electric gain measurements and the optical gain above.
* Using a mixer and SR560(s), make a separate feedback circuit for the PZT lock. Then use the PC path
  to measure the PC path response.
* See the response of the FSS board to large impulse/step inputs to find the cause of the PC path craziness.
etc ...
Attachment 1: OPLTFs.pdf
OPLTFs.pdf
Attachment 2: OPLTFsScaled.pdf
OPLTFsScaled.pdf
  904   Fri Aug 29 18:24:48 2008 ranaHowToPSLPMC: PZT Calibration
I calibrated the PMC PZT at DC by using 'trianglewave' to drive the DC offset slider
and reading back PMC_PZT and PMC_TRANSPD_F (both are DC coupled DAQ channels).

The attached PDF illustrates the method: look at the voltage required to span 1 FSR and then divide.
PMC_cal (m/V) = (1064 nm)/2 / V_FSR
The calibration for our PZT is therefore 10.4 nm/V.
The full scale (0-300 V) range is 3.1 microns.

From Jenne's elog entry we know that the series resistor to the PZT is 63.6 kOhms. The PZT is labeled as
having a capacitance of 279 nF. So the PMC drive's pole frequency is 1/2/pi/63.6e3/279e-9 = 9 Hz +/- 0.5 Hz.
The cable capacitance is ~20 pF/foot so its not significant for this.

The template file is Templates/PMC-PZTcal.xml.

Using the above calibrations, also plot the calibrated PMC ERR and PZT spectra.
Attachment 1: pmc-pzt-cal.pdf
pmc-pzt-cal.pdf
Attachment 2: mcf.png
mcf.png
  903   Fri Aug 29 17:39:25 2008 ranaConfigurationPSLPMC: ADC Channels
The attached PNG shows the PMC error and controls signals with no calibration.

There are 3 states:

DARK - RF input disabled & output blanked. This should be a measure of the ADC noise

(-10 dB) - This is with the gain slider down at 5 dB instead of the nominal 15 dB.

Looks like the Generic DAQ board whitening is good enough for these signal levels above ~1 Hz.

From the low and high gain spectra it also looks like the UGF is ~500 Hz with the gain at 15 dB.
Attachment 1: mcf.png
mcf.png
  902   Fri Aug 29 16:35:18 2008 YoichiConfigurationPSLbeam block distorted

Quote:
There was a beam block after the Mach Zender. Who or what put this there?

The going to the MC now looks distorted as if someone has left something funny in the beam or maybe the new PMC has started to degrade??

Use the ELOG people...its good for you.


The apparent distortion of the MC refl. was caused by mis-alignment of the MC mirrors.
Because the MC1 was mis-aligned, the reflected light was clipped by a steering mirror.
I restored the MC angle bias values from the conlog history and now the MC locks.
According to conlog, the MC alignment was changed at around 18:30 on Thursday PDT.
It could have been caused by the computer reboots.
  901   Fri Aug 29 15:01:45 2008 steveUpdatePSLMOPA_HTEMP in increasing
The laser chiller temp is 21.9C ( it should be 20.0C )
Control room temp 73F ok, no obvious block

Ops, there is a piece of paper blocking the intake of the chiller

This is a four day plot. The paper was blocking the air flow all day.
Attachment 1: htcl.jpg
htcl.jpg
  900   Fri Aug 29 12:43:44 2008 josephbSummaryComputersc1susvme1 down
Around noon today, c1susvme was having problems. The C0DAQ_RFMNETWORK light was red. The status light was off, the sig det light was amber and the own data light was green. I could also ssh in, but could not not run startup. I switched off the watchdogs for c1susvme2 (the watchdogs for c1susvme1 had already been tripped), and manually power cycled the crate.

However, when c1susvme1 when it came back up it had not mounted the usual cvs/cds/ directories. c1susvme2 did however. c1susvme1 has been on the new network for awhile, while c1susvme2 was switch over today. So apparently switching networks doesn't help this particular problem.

I did a remote reboot of c1susvme1, and it came up with the correct files mounted. Both machines ran their approriate startup.cmd files and are currently green.
  899   Fri Aug 29 12:41:26 2008 josephb, EricConfigurationComputersMore front ends moved to new network
Used Cat6 cables to finish moving all the front ends in 1Y4 and 1Y5 over to the new GigE network switches, specifically to the switch in 1Y6. This included the ones labeled c1susvme2, c1sosvme, and c1dscl1epics0.
  898   Fri Aug 29 11:05:11 2008 josephbSummaryComputersc1asc was down this morning
I had to manually reboot c1asc this morning, as for some unknown reason its status was red, and the fiber lights on the board were status:red, sig det:amber, own data: nothing. Shut the crate down, turned it back on, heard a beep, then followed wiki reboot instructions. Seems to be working now.
  897   Fri Aug 29 11:01:49 2008 josephbConfigurationComputersAttempt to change a channel gain in ICS-110B
As noted earlier by Rana, I was playing around with the /cvs/cds/caltech/chans/daq/C1IOOF.ini file with help from Rob. I had made a backup before hand and saved it as C1IOOF.ini.Aug-28-2008. (I have since been informed that C1IOOF.ini.082808 would have been prefered as a name).

We had been trying to up the gain in the C1: PSL-ISS_INMONPD_F in order to do a very low power PMC sweep, in an attempt to get clean modes for fitting. Initially we pressed the reconfig button on the C0DAQ_DETAIL screen, but all that seemed to do was change the Config File CRC. We proceeded to reboot fb40m remotely. However, any change to the ini file (even an extra space at the end of the file) caused a 0x2000 status for C1IOVME16k on the C0DAQ_DETAIL screen. At the time I presumed it was comparing the CRC of the ini-file to something else.

Digging around on in Alex's webspace at http://www.ligo.caltech.edu/~aivanov/ , I found the NDS Access page, which indicated that 0x2000 was a conflict between the front-end and frame builder .ini files.

"There is also status bit 0x2000 which gets added when the DCU configuration is different in front-end and frame builder. That is you can change and .ini file an then reload DAQ configuration with Epics button, which reconfigures the front-end, but leaves frame builders with invalid old configuration. They will detect this change and set the status to 0x2000 to indicate this condition. You will have to restart frame builders to pick up new .ini file and set status back to zero for the affected DCU."

It was when I was going to try reseting the c1iovme via the C0DAQ_RFMNETOWRK medm screen that we realized the EPICS controls were not responding properly. The .ini file was returned to its original form, and mass reboots commenced.
  896   Fri Aug 29 10:20:32 2008 YoichiConfigurationPSLbeam block distorted

Quote:
There was a beam block after the Mach Zender. Who or what put this there?

The going to the MC now looks distorted as if someone has left something funny in the beam or maybe the new PMC has started to degrade??

Use the ELOG people...its good for you.


I put the block. I was frequently reaching to the FSS box to change the test point probes. I put the block to protect my hands/clothes from being burnt accidentally.
  895   Fri Aug 29 02:40:43 2008 rana,jenneUpdatePSLPMC Servo Board

Quote:
Board is back in. PMC is locked.


This entry has details about the low pass filter after the PMC mixer. This filter has a few purposes:

1] Remove the beat signal (at 2*f_mod) between the PD RF signal at f_mod and the LO signal at f_mod.
2] Remove the beat signal (at f_mod) between the PD RF signal at 2*f_mod (which comes from the
beating of the upper and lower RF sidebands) and the LO signal at f_mod.
3] Remove other RF signals from non-ideal behavior of the LO drive signal and distortion in the RF PD pre-amp.


So its important to have a very good rejection at 35 MHz and higher. I used the Hartmut LC network design which is
installed on H1, H2, & L1. Since there is a high gain in the audio amps right after the mixer we have to get rid of
the RF or else we'll get slew rate limited or otherwise rectified downconversion of the RF signal into our audio band.

Of course, what everyone immediately realizes from the above 3 points, is that this filter can't protect the PMC
noise performance from homodyne mixing (e.g. 2*f_mod in the LO and 2*f_mod in the RF PD). To get around that, we're
ordering some filters from Mini-Circuits to remove the 2f from those signals by ~30 dB. As long as we install
the same filters on the RF and LO legs, there should be no significant phase shift in the demodulated signal.

The attached 2 page PDF shows the calculated before and after TFs of this filter. The 2 attached .m files
calculate the TF's and have ascii art which shows how the filter works.

Here's a comparison of the attenuation (in dB) of 2 candidate Mini-circuits filters:

f(MHz)SLP-30SLP-50
31 0.5 0.4
35 1.3 0.4
38 6.1 0.4
40 10.8 0.42
61 46.3 14.8
71 60 29
91 76.9 48
10780 60

We don't have tabulated data at the same frequencies for both filters so I just made up some of the points by eye-balling the
plots from the catalog - but you get the idea: we can get away with using the SLP-30 at 35 MHz since it only attenuates the
signals by ~1.5 dB. So if someone can find 4 of these then Steve doesn't have to order any from Mini-Circuits.
Attachment 1: pmclp-07.pdf
pmclp-07.pdf pmclp-07.pdf
Attachment 2: pmclp_40m_080824.m
% PMCLP is a TF of the IF filter after the PMC mixer
%       

% Mixer_Voltage -- Rs -- L1 --- L2 ---------Vout
%                            |      |   |
%                           C1     C2   Rl                   
%                            |      |   |
%                           GND    GND  GND
%

... 58 more lines ...
Attachment 3: pmclp.m
% PMCLP is a TF of the IF filter after the PMC mixer
%       

% Mixer_Voltage -- Rs -- L1 --- L2 ---------Vout
%                            |      |   |
%                           C1     C2   Rl                   
%                            |      |   |
%                           GND    GND  GND
%

... 57 more lines ...
  894   Thu Aug 28 19:02:25 2008 rana, josephb, robSummaryComputersbig boot
This afternoon Joe did something with an .ini file (look for his detailed elog entry) and the computers went bad.
RFM network screen not active - filter modules not working.

We went around and booted every machine as has been done before. The correct order for a memory corruption
fixing big boot is the following:

    [1] RESET the RFM switches near the FB racks.
    [2] Power cycle c1dcuepics.
    [3] Power cycle all other crates with real time CPUs:
    c1iscey, daqctrl, daqawg, c1susvme1, c1susvme2, c1sosvme, c1iovme, c1lsc, c1asc, & c1iscex
    [4] Start up all FEs as described in Wiki.
    [5] Burt restore everyone (losepics, iscepics, assepics, omcepics?)
  893   Thu Aug 28 18:56:14 2008 ranaConfigurationPSLbeam block distorted
There was a beam block after the Mach Zender. Who or what put this there?

The going to the MC now looks distorted as if someone has left something funny in the beam or maybe the new PMC has started to degrade??

Use the ELOG people...its good for you.
  892   Wed Aug 27 13:55:43 2008 rana,jenneUpdatePSLPMC Servo Board
Board is back in. PMC is locked.

Nominal gain is now 15 dB with brick. We need to do more studies:

  • Find out why there is still 35 MHz signal at the error point. Order some low pass filters to cut off above 35 MHz.
  • Explore brick + no-brick loop shapes and error spectra.
  • Measure and set the OLG.

We've left the copper-wrapped lead brick installed to let it slowly conform to the glass better.
  891   Wed Aug 27 12:09:10 2008 EricSummaryCamerasWeekly Summary
I added a configuration file parser to the Snap code. This allows all command line parameters (like exposure time, etc.) to be saved in a file and loaded automatically. It also provides a method of loading parameters to transform a point from its location on the image to its location in actual space (loading these parameters on the command line would substantially clutter it). The code is now fully set-up to test servo-ing one of the mirrors again, and I will test this as soon as the PMC board stops being broken and I can lock the X-arm.

I also took an image of the OSEMs on ETMX in order to apply the rotation transform code in order to determine the parameters to pass to Snap. The results were alpha = 2.9505, beta = 0.0800, gamma = -2.4282, c = 0.4790. These results are reasonable but far from perfect. One of the biggest causes of error was in locating the OSEMs: it is difficult to determine where in the spot of light the OSEM actually is, and in one case, the center was hidden behind another piece of equipment. Nevertheless, the parameters are good enough to use in a test of the ability to servo, though it would probably be worth trying to improve them before using them for other purposes. The original and rotated images are attached.

I've begun working on calculations to figure out how much power loss can occur due to a given cavity misalignment or change in a mirror's radius of curvature from heating. The goal is to determine how well a camera can indirectly detect these power losses, since a misalignment produces a change in beam position and a change in radius of curvature produces a change in beam waist, both of which can be measured by the camera.

Joe and I hunted down the requisite equipment to amplify the photodiode at the output of the PMC, allowing us to turn the laser power down even more during a scan of the PMC, hopefully avoiding thermal effects. This measurement can be done once the PMC works again.
Attachment 1: originalETMX.png
originalETMX.png
Attachment 2: rotatedETMX.png
rotatedETMX.png
  890   Wed Aug 27 10:55:35 2008 YoichiHowToComputersAnnoying behavior of the touch pads of the lab. laptops is fixed
I was sick of the stupid touch pad behavior of the lab. laptops, i.e. firefox goes back and forth in the history when the cursor is moved.
It was caused by firefox mis-interpreting the horizontal scroll signal as back/forward command.
I stopped it by going to about:config in firefox and set mousewheel.horizscroll.withnokey.action to 0 and
mousewheel.horizscroll.withnokey.sysnumlines to true.
  889   Tue Aug 26 19:07:37 2008 YoichiHowToComputersReading data from Agilent 4395A analyzer through GPIB from *Linux* machine
I succeeded in reading data from Agilent 4395A analyzer, who's floppy is crappy, through GPIB from a Linux machine using
agilent 82357B USB-GPIB interface.
I installed the linux GPIB driver to one of the lab. laptops (the silver DELL one currently sitting on the 4395A analyzer).
I wrote an initialization script for the USB-GPIB interface and a small python script for reading data from the analyzer.

[Usage]

1. Connect the USB-GPIB interface to the laptop and the analyzer.
2. Run /usr/local/bin/initGPIB command (it takes about 10sec to complete).
3. Run /usr/local/bin/getgpibdata.py > data.txt to save data from the analyzer to a text file.

The data format is explained in the comments of getgpibdata.py
This method is way faster than the unreliable floppy. The data is transfered in a few sec.

I'm now writing a wiki page on this
http://lhocds.ligo-wa.caltech.edu:8000/40m/GPIB

I will install the same thing into the other DELL laptop soon.
Let me know if you have trouble with this.
  888   Tue Aug 26 18:19:16 2008 ranaOmnistructureElectronicsResistor Noise at the 40m
As Stefan points out in his recent ISS ilog entries at LLO, Daniel Sigg recently wrote a
recommendation memo on resistor and capacitor
choices: T070016.

While working on the PMC I have had to use leaded resistors and wondered about the noise. As it turns
out we have the RN series of 1/4 W resistors from Stackpole Electronics. The RN series are
metal film resistors (datasheet attached); metal film is what Sigg recommends for lowest flicker
noise.

So we are OK for using the Stackpole 1/4 W leaded resistors in low noise circuits.
Attachment 1: SEI-RN_RNM.pdf
SEI-RN_RNM.pdf SEI-RN_RNM.pdf
  887   Tue Aug 26 15:06:16 2008 steveConfigurationVACrga scan
Pumpdown 66 PRM-maglev vac normal -day 11
short form: pd66PRM-m-d11
Attachment 1: RGA-0808260125.png
RGA-0808260125.png
  886   Tue Aug 26 12:00:45 2008 JenneSummaryPEMTransfer function of Ranger seismometer
This finishes up the calibration that Rana started in elog # 881.

The calibration of the Ranger seismometer should also include:
2 zeros at 0 Hz
2 poles at 1.02 Hz

This comes from finding the transfer function between the mass's motion and the motion of the ground.
    ..
m * x  = (x_G - x) * k  + d(x_G - x) * b
                          dt

where
  • m = mass
  • x = displacement of the mass
  • x_G = displacement of the ground
  • k = spring constant
  • b = damping constant

This gives
x               w0^2  +  i*w*w0/Q
----    =    -----------------------
x_G           w0^2 + i*w*w0/Q - w^2

where
  • w0 = sqrt(k/m) = natural frequency of spring + mass
  • w = frequency of ground motion
  • Q = q-factor of spring + mass system = 1/2 for critically damped system

The readout of the system is proportional to
d  (x - x_G)          (    w0^2  +  i*w*w0/Q          )    .                    w^2               .
dt                 =  (  -----------------------  - 1 ) * x_G   =      ----------------------- * x_G
                      (   w0^2 + i*w*w0/Q - w^2       )                w0^2 + i*w*w0/Q - w^2
Since we read out the signal that is proportional to velocity, this is precisely the transfer function we're looking for. With w0 = 1.02 Hz and Q = 1/2 for the critically damped system, we have 2 zeros at 0 and 2 poles at 1.02.
  885   Tue Aug 26 09:58:59 2008 steveOmnistructureCOCETMX is #03
This is the picture of ETMX from the upper south west viewport
Attachment 1: etmx-d40posg_012.jpg
etmx-d40posg_012.jpg
Attachment 2: etmx03.jpg
etmx03.jpg
  884   Tue Aug 26 09:04:59 2008 ranaConfigurationPSLPMC Servo Board: Out for Repairs
I've started modifying our PMC board to bring it up to the 21st century - leave the screen alone or else you might zap something.
  883   Mon Aug 25 21:15:23 2008 ranaConfigurationLSCaux NPRO off
Looks like no one has used the Lightwave NPRO on the AS table after Koji left, so I turned it off so that it can rest until Alberto does the X-arm measurements.
  882   Mon Aug 25 17:45:34 2008 rana, josephb, robHowToPEMAccelerometer range
Joe shows us by jumping up ~15" in the control rom that the accelerometers are set with not enough gain.

Since this is taken around 5:30 in the evening, so we can take the nearby time series to represent what a
high noise level is. I recommend we up the gain using the ICS-110B .ini file.
Attachment 1: Screenshot-4.png
Screenshot-4.png
  881   Mon Aug 25 15:50:18 2008 ranaSummaryPEMRanger SS-1
The manual for the Ranger SS-1 seismometer can be found on line here:
ftp://ftp.kmi.com/pub/software_manuals/300190/300190nc.pdf

and now in our 40m PEM Wiki page:
Ranger_SS-1

To calibrate it, we use the formula from the manual:
                 R_x
G_L = G_0 * ------------   =  149 +/- 3 V/(m/s)
             R_x  +  R_c

where
G_0 = 340 V/(m/s)    (generator constant)
R_x = 4300 Ohms      (external damping resistor in Pomona box)
R_c = 5500 Ohms      (internal coil resistance)

Then we have a gain of 200 in the SR560 so that gets us to ~30000 V/(m/s).

And then there's a DAQ conversion factor of the usual 2^16 cts / 4 V.

so the calibration constant is

G = 488 counts / (micron/sec)

in the ~1-50 Hz band
  880   Mon Aug 25 14:42:09 2008 EricConfigurationCamerasETMX Digital Camera
I changed the lens on the camera looking at the ETMX to a 16mm, 1:1.4 zoom lens. This is in preparation to measure a couple parameters that depend on the camera's position and angle, so please avoid repositioning it for a couple of days.
ELOG V3.1.3-